OpenStudy (anonymous):

solve the following quadratic equation by the square root property of equality x^2-81=0

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

(9,-9) ?

OpenStudy (anonymous):

x^2=81=>x^2=(9)^2=>x=+-9